OpenStudy (anonymous):

if 1 is root of squareroot of a-x = square root of b=x=squareroot of a=b=2012 then possible values of a,b

OpenStudy (anonymous):

umm, question is a bit vague. can u come over again?

OpenStudy (akashdeepdeb):

A bit vague? It doesn't make any sense at all!

OpenStudy (akashdeepdeb):

\[\sqrt{a-x} = \sqrt{b} = x = \sqrt{a} = b = 2012\] ?